A simple pendulum that consist of a rock of mass 3.5 kg tied to the end of a (massless) string of

length 1.5 m.

(a) What is the frequency of the pendulum?

(b) What is the period of the pendulum?


1
Expert's answer
2021-01-15T13:47:07-0500

a)


"f=\frac{1}{2\pi}\sqrt{\frac{g}{l}}\\f=\frac{1}{2\pi}\sqrt{\frac{9.8}{1.5}}=0.41\ Hz"

b)


"T=2\pi\sqrt{\frac{l}{g}}\\T=2\pi\sqrt{\frac{1.5}{9.8}}=2.5\ s"